Find the derivative of y= sinx/2+cotx


1
Expert's answer
2021-02-24T06:01:55-0500

Find the derivative of "y=\sin{\frac{x}{2}}+\cot{x}"

Solution:

"y'=\frac12\cos{\frac {x}{2}}-\frac{1}{\sin^2{x}}"

Answer:

"y'=\cos{\frac {x}{2}}\cdot(\frac x2)'-\frac{1}{\sin^2{x}}=\frac12\cos{\frac {x}{2}}-\frac{1}{\sin^2{x}}" .